OpenStudy (pottersheep):

Why can't 1/squareroot x = a negative number?

OpenStudy (pottersheep):

please please help

OpenStudy (across):

\[\frac{1}{\sqrt{x}}\not<0\]

OpenStudy (across):

Is that what you're asking?

OpenStudy (pottersheep):

not really :s

OpenStudy (pottersheep):

thnx anyways

OpenStudy (anonymous):

Well, in a sense, it can be a negative number. For example, if x is 4, then it could be:\[\frac{1}{\sqrt{4}} = \frac{1}{-2}\]Or\[\frac{1}{\sqrt{4}} = \frac{1}{2}\]

jimthompson5910 (jim_thompson5910):

The square root of any nonnegative number is going to be a nonnegative number. In other words, the square root of some number will NEVER be negative. So sqrt(x) is either 0 or it's positive So 1/sqrt(x) is ALWAYS positive because you can't divide by zero

OpenStudy (anonymous):

I think the point pottersheep is getting at is the idea that for 1/sqrt(x), the x can't be negative otherwise you will return an error or and imaginary number, especially handy to consider when studying domains and ranges of functions